Solve the equation :


(7y-3x+3)dy + (3y-7x+7)dx =0

1
Expert's answer
2022-01-24T02:49:00-0500

(7y-3x+3)dy + (3y-7x+7)dx =0

(3x-7y-3)dy=(3y-7x+7)dx

"\frac{dy}{dx}=\frac{3y-7x+7}{3x-7y-3}\\
now,let ~x=x'+h~and~y=y'+k\\
\frac{dy'}{dx'}=\frac{3y'-7x'+3k-7h+7}{3x'-7y'+3h-7k-3}\\
putting~h=1~and~k=0,we~get\\
\frac{dy'}{dx'}=\frac{3y'-7x'}{3x'-7y'}\\
now,let~ y'=vx'~so~that\\
\frac{dy'}{dx'}=v+x'\frac{dv}{dx'}\\
we~have\\
v+x'\frac{dv}{dx'}=\frac{3vx'-7x'}{3x'-7vx'}\\
x'\frac{dv}{dx'}=\frac{3v-7}{3-7v}-v\\
x'\frac{dv}{dx'}=\frac{3v-7-3v+7v^2}{3-7v}\\
x'\frac{dv}{dx'}=\frac{7v^2-7}{3-7v}\\
\frac{dx'}{x'}=\frac{3-7v}{7v^2-7}dv\\
integrating~both~sides\\
\int \frac{dx'}{x'}=\int \frac{3}{7v^2-7}dv-\int \frac{7v}{7v^2-7}dv\\
ln|x'|+c= \frac{3}{14}ln \frac{|v-1|}{|v+1|}-\frac{1}{2}ln|{v^2-1}|\\
ln|x'|+c= \frac{3}{14}ln \frac{|y'-x'|}{|y'+x'|}-\frac{1}{2} ln|\frac{(y')^2-(x')^2}{(x')^2}|\\
Recall~v=\frac{y'}{x'}\\
ln|x'|+c= \frac{3}{14}ln {|y'-x'|}- \frac{3}{14}{|y'+x'|}-\frac{1}{2} ln|{y'-x'}|+\frac{1}{2} ln|{y'+x'}+\frac{1}{2} ln|({x'})^2|\\
ln|x'|+c= -\frac{2}{7}ln {|y'-x'|}-\frac{5}{7}ln {|y'+x'|}+\frac{1}{2} ln|({x'})^2|\\
2ln|y'-x'|+5ln|y'+x'|=C\\
2ln|y-x+1|+5ln|y+x-1|=C\\ 

(Since~y'=y~and~x'=(x-1))"